OpenStudy (anonymous):

What is the value of 1*1! + 2*2! + 3!*3! + ............ n*n! ?

OpenStudy (aravindg):

nikvist can u help me??

OpenStudy (nikvist):

\[S_n=\sum\limits_{k=1}^{n}k\cdot k!=\sum\limits_{k=1}^{n}(k+1-1)\cdot k!=\sum\limits_{k=1}^{n}(k+1)!-\sum\limits_{k=1}^{n}k!=\]\[=(n+1)!-1\]

OpenStudy (nikvist):

ok, AravindG !!! I will help you

OpenStudy (aravindg):

:)